OpenStudy (anonymous):

You bought a sports car for $89,000. What will the car be worth after 10 years if the depreciation rate is 8%? A. $38,660.57 B. $50,339.43 C. $82,130.93 D. $204,885.74

OpenStudy (anonymous):

A I believe...

OpenStudy (anonymous):

It will be A

OpenStudy (anonymous):

of course it wont be C OR D. The calculation is extremely tedious actually. But still; rate = 8% , and Principal = 89000, time (n) = 10 years. We can use the normal compound interest formula A = P(1-R/100)^n (1-r/100 since its depreciation) A = 89000(1-8/100)^10 => 89000(1-2/25)^10 => 89000(25-2/25)^10 => 89000(23/25)^10 which on solving(calculator) gives $38660.58

OpenStudy (anonymous):

anyway i dont have 89000$ as of now to buy any car. :P